Need help finding the ho on this quadratic word problem the answer I tried to figure out is 6.74 seconds. The word problem is For an object dropped from a height ho above the ground,its height after t seconds is given by h(t)=-16t^2+ho,where H is measured in feet. A ball is dropped from the top 96 foot tall building. How long will it take to fall to the ground? How long will it take to fall half of the distance to ground level?
Im not really good with word problems

Answer by lwsshak3(11628)

The word problem is For an object dropped from a height ho above the ground,its height after t seconds is given by h(t)=-16t^2+ho,where H is measured in feet. A ball is dropped from the top 96 foot tall building. How long will it take to fall to the ground? How long will it take to fall half of the distance to ground level?
**
h(t)=-16t^2+ho
For given problem, ho=96
h(t)=-16t^2+96
..
When object falls to the ground, h(t)=0
0=-16t^2+96
16t^2=96
t^2=96/16=6
t=√6
..
When object falls half the distance to the ground, h(t)=48
48=-16t^2+96
16t^2=48
t^2=48/16=3
t=√3
ans:
When object falls to the ground, it will take√6 seconds
When object falls half the distance to the ground,it will take √3 seconds